# Source-Channel Separation Theorems for Distortion Perception Coding

### v1

@@ Abstract
It is well known that separation between lossy source coding and channel coding is asymptotically optimal under classical additive distortion measures. Recently, coding under a new class of quality considerations, often referred to as perception or realism, has attracted significant attention due to its close connection to neural generative models and semantic communications. In this work, we revisit source-channel separation under the consideration of distortion-perception. We show that when the perception quality is measured on the block level, i.e., in the strong-sense, the optimality of separation still holds when common randomness is shared between the encoder and the decoder; however, separation is no longer optimal when such common randomness is not available. In contrast, when the perception quality is the average per-symbol measure, i.e., in the weak-sense, the optimality of separation holds regardless of the availability of common randomness.
@@ I Introduction
Recently, a new type of quality consideration, often referred to as perception or realism, has emerged [4]. Roughly speaking, the requirement is that the reconstruction should have a probability distribution that is similar to the original source distribution. This consideration is well-motivated in neural compression, as neural generative models enjoy strong performance with visually pleasing reconstructions, which may however be viewed as having low quality under the traditional distortion measures such as mean squared error (MSE). Moreover, semantic communication systems may also need to be measured under such a quality consideration. Previous efforts in the study of coding with perception quality consideration have focused on the source coding formulation, i.e., the rate-distortion-perception function. Several conclusive results have been reported [5, 6, 7, 8, 9].
It should be noted that the perception quality consideration is different from the classical single-letter additive distortion measure, since the quality may be measured on the whole coding block instead of on individual sample positions. Therefore, the question of whether separation is optimal under the perception quality consideration remains mostly open. In this work, we revisit source-channel separation under the distortion-perception consideration. We show that when the perception quality is measured on the block level, i.e., in the strong-sense, the optimality of separation still holds when common randomness is shared between the encoder and the decoder; however, separation is no longer optimal in general when such common randomness is not available. In contrast, when the perception quality is the average per-symbol measure, i.e., in the weak-sense, the optimality of separation holds regardless of the availability of the common randomness.
Our result for the case under strong-sense perception measure with common randomness is particularly notable, since the proof does not rely on a single-letter characterization of the corresponding rate-distortion-perception region. This proof allows us to establish the separation result under a general set of conditions.
@@ II-A Notation and Preliminaries
The channel we consider in this work is the classic discrete memoryless channel given by $P_{Y|X}$, and we denote its capacity as $C$. We assume the discrete memoryless source follows a probability distribution $P_{S}$ on a finite alphabet $\mathcal{S}$. A distortion measure function $\delta:\mathcal{S}\times\mathcal{S}\rightarrow[0,\infty)$ and a sequence of perception measure functions $d_{n}:\Delta_{\mathcal{S}^{n}}\times\Delta_{\mathcal{S}^{n}}\rightarrow[0,\infty]$ are assumed, where $\Delta$ is the probability simplex on the alphabet specified by the subscript.
@@
We state our assumptions on the perception measure below.
Assumption 1.
The normalized perception measure $\frac{1}{n}d_{n}(\cdot,\cdot)$ is bounded.
@@ II-B DP Joint Source-Channel Coding Achievable Regions
A distortion and perception pair $(D,P)$ is called achievable on a channel with a mismatch factor $\kappa$, if for all sufficiently large $n$, there exist common randomness $U$, an encoding function $f:\mathcal{S}^{n}\times\mathcal{U}\rightarrow\mathcal{X}^{m}$, a decoding function $g:\mathcal{Y}^{m}\times\mathcal{U}\rightarrow\mathcal{S}^{n}$, such that $\hat{S}^{n}=g(Y^{m},U)$ is the reconstruction and $Y^{m}$ is the channel output on the memoryless channel $P_{Y|X}$ with input vector $X^{m}=f(S^{n},U)$, which satisfy the condition that
@@
The first condition above specifies that the source-channel mismatch factor must be upper-bounded properly, and the second condition and the last condition specify the distortion constraint and the (strong-sense) perception constraint, respectively. There are other related concepts in the literature that are also of practical interests. We consider the following cases in our work:
@@ II-C RDP Coding Achievable Regions
Similar to the joint coding settings, we can define the weak-sense rate-distortion-perception achievable region, the strong-sense rate-distortion-perception achievable region without common randomness, and the weak-sense rate-distortion-perception achievable region without common randomness, respectively. We denote these regions as $\mathcal{G}_{w}$, $\hat{\mathcal{G}}$, $\hat{\mathcal{G}}_{w}$, respectively.
@@
Under the assumptions above, the distortion-perception joint coding achievable region $\mathcal{J}$ and the rate-distortion-perception achievable region $\mathcal{G}$ are both convex, regardless of whether the strong-sense or the weak-sense perception measure is adopted.
@@ II-D DP Separate Coding Regions
We can similar define the weak-sense separate coding achievable distortion-perception region, the strong-sense separate coding achievable distortion-perception region without common randomness, and the weak-sense separate coding achievable distortion-perception region without common randomness, denoted as $\mathcal{J}^{\prime}_{w}$, $\hat{\mathcal{J}}^{\prime}$, and $\hat{\mathcal{J}}^{\prime}_{w}$, respectively.
We also need the following rate-distortion-perception function in the proof:
$R(D,P)=\min_{(R,D,P)\in\mathcal{G}}R.$ (10)
@@
III Coding under the Strong-Sense Perception Measure
@@ III-A Separation: Common Randomness Present
Under Assumption 1 and 2, we have $\mathcal{J}^{\prime}\supseteq\mathcal{J}$.
A further technical assumption is needed for the forward direction.
Assumption 3.
@@
Under Assumptions 1, 2, and 3, we have $\mathcal{J}^{\prime}\subseteq\mathcal{J}$.
As a consequence, under Assumptions 1, 2, and 3, we have $\mathcal{J}^{\prime}=\mathcal{J}$, which essentially says separate coding achievable region is exactly the same as joint coding achievable region under these conditions. We remark that a single-letter characterization of the achievable rate-distortion-perception region is not known under Assumptions 1, 2, and 3. The single-letter results given in [7] were in fact established under a more stringent requirement on the perception functions. Nevertheless, we are still able to establish the optimality of the separation architecture using a proof that does not rely on the single-letter characterization.
@@ III-B Failure of Separation: Common Randomness Absent
Under the strong-sense perception requirement, when there does not exist common randomness shared by the encoder and the decoder, source-channel separation is not optimal in general. We next provide a counter-example from [10] to illustrate this.
@@ IV Coding under the Weak-Sense Perception Requirement
Under Assumptions 1 and 3, $\mathcal{J}_{w}=\mathcal{J}^{\prime}_{w}=\hat{\mathcal{J}}_{w}=\hat{\mathcal{J}}^{\prime}_{w}=\mathcal{J}^{*}_{w}$, where
@@ V-A Proof of Theorem 1
We wish to show that there exists rate-distortion-perception code that achieves with $R\leq\kappa C+\delta$ for any $\delta>0$. By the strong functional representation lemma [11], there exists a random variable $U^{\prime}$, independent of $S^{n}$ such that $\hat{S}^{n}$ can be expressed as a deterministic function of $(S^{n},U^{\prime})$ and
@@
for some positive $\delta_{n,m}$ which diminishes as $n,m$ becomes large.
Since our new code preserves $p_{S^{n},\hat{S}^{n}}$, we have
@@ V-B Proof of Theorem 2
$\{(D,P):(R,D,P)\in\mathcal{G},R<\kappa C\},$ (19)
as the rate-distortion-perception function is continuous. The continuity is implied by the fact that the joint coding achievable distortion-perception region is closed by definition, and the rate-distortion-perception function is continuous in the relative interior of the non-negative quadrant (implied by its convexity).
@@
With this code, let us consider its length-$k$ extension, with $U_{i}$’s mutually independent for $i=1,2,\ldots,k$. Let $P_{S^{nk},\hat{S}^{nk}}$ denote the joint distribution of $(S^{nk},\hat{S}^{nk})$ induced by this extended code.
When $k$ is large, we have $\frac{1}{nk}H(Z^{k}|U^{k})\leq R+\delta$ holds with high probability. More precisely, choose an arbitrary pair $\delta_{R},\epsilon_{R}>0$, there exists a sufficiently large $k$ such that
$Pr\left(\text{the total description length of $Z^{k}$}\geq nk(R+\epsilon+\delta_{R})\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq\epsilon_{R}.$ (23)
We shall send this description of the $k$-extension of the original rate-distortion perception code, if the total description length is less than $nk(R+\epsilon+\delta_{R})$, using an almost capacity-achieving channel code; otherwise, we simply send a fixed dummy sequence of a shorter length (correspondingly it will be viewed as an error event).
Since $R<\kappa C$, when $n,k$ are both sufficiently large, there exists a channel code at rate greater than $\kappa^{-1}(R+\epsilon+\delta_{R}+\frac{1}{nk})$, such that the decoding error probability is upper-bounded by $\epsilon_{C}$, where $\epsilon_{C}$ diminishes as $n,k$ are sufficiently large. The new decoder first attempts to recover the intended $Z^{k}$ sequences, and let $\hat{Z}^{k}$ denote this reconstruction. Two error events (when $Z^{k}\neq\hat{Z}^{k}$) may occur: 1) The description of the concatenated source sequence $Z^{k}$ is too long and the fixed dummy sequence was sent, and 2) A channel decoder error occurred. By the union bound, we have
@@
$\leq\frac{1}{nk}\sum_{t=1}^{nk}\mathbb{E}_{P}[\delta(S_{t},\hat{S}_{t})]+(\epsilon_{R}+\epsilon_{C})\delta_{\max}$
$\leq D+\epsilon+(\epsilon_{R}+\epsilon_{C})\delta_{\max},$ (26)
and
@@
where (27) is due to the fact that $d_{nk}$ is continuous with respect to $d_{TV}$ and sub-decomposable. ∎
@@ V-C Proof of Theorem 3
In the weak-sense setting, it is clear by definition that $\mathcal{J}_{w}\supseteq\hat{\mathcal{J}}_{w}$, $\mathcal{J}_{w}\supseteq\mathcal{J}^{\prime}_{w}$, and $\hat{\mathcal{J}}_{w}\supseteq\hat{\mathcal{J}}^{\prime}_{w}$. It was also shown in [7] that the weak-sense achievable rate-distortion-perception region is characterized as
@@
To show $\mathcal{J}_{w}\subseteq\mathcal{J}^{*}_{w}$, consider an arbitrary pair $(D,P)\in\mathcal{J}_{w}$. For any $\epsilon>0$, there exists a perception-distortion code $(f,g)$ that achieves $(D+\epsilon,P+\epsilon)$. For this code we have
@@
$P+\epsilon\geq\frac{1}{n}\sum_{t}\mathbb{E}d_{1}(S_{t},\hat{S}_{t})=\mathbb{E}d_{1}(S,\hat{S}).$ (33)
Since $\epsilon$ can be made arbitrarily small by making $n$ large, we have $(D,P)\in\mathcal{J}^{*}_{w}$.
To show $\hat{\mathcal{J}}^{\prime}_{w}\supseteq\mathcal{J}^{*}_{w}$, we shall rely on the combination of the coding scheme given in [7] for the rate-distortion-perception problem and the conventional capacity-achieve channel codes. Similarly to the proof of Theorem 2, the only issue that needs to be resolved is that due to excess source coding description length and channel uncertainty, the description $Z$ has a diminishing error probability of being decoded erroneously at the decoder. However, the penalties such induced on the distortion and the perception are both negligible, when Assumptions 1 and 3 hold. The proof is thus complete ∎
@@ VI Conclusion
We consider source-channel separation with perception consideration in addition to the standard distortion measure. We show that under the strong-sense perception constraint, separation is optimal when common randomness is available at the encoder and the decoder, however, when common randomness is not available, source-channel separation is not optimal in general. In contrast, under the weak-sense perception constraint, separation is always optimal regardless of the existence of common randomness.

### v2

@@ Abstract
It is well known that separation between lossy source coding and channel coding is asymptotically optimal under classical additive distortion measures. Recently, coding under a new class of quality considerations, often referred to as perception or realism, has attracted significant attention due to its close connection to neural generative models and semantic communications. In this work, we revisit source-channel separation under the consideration of distortion-perception. We show that when the perception quality is measured on the block level, i.e., in the strong sense, the optimality of separation still holds when common randomness is shared between the encoder and the decoder; however, separation is no longer optimal when such common randomness is not available. In contrast, when the perception quality is the average per-symbol measure, i.e., in the weak sense, the optimality of separation holds regardless of the availability of common randomness.
@@ I Introduction
Recently, a new type of quality consideration, often referred to as perception or realism, has emerged [4]. Roughly speaking, the requirement is that the reconstruction should have a probability distribution that is similar to the original source distribution. This consideration is well-motivated in neural compression, as neural generative models enjoy strong performance with visually pleasing reconstructions, which may, however, be viewed as having low quality under the traditional distortion measures such as mean squared error (MSE). Moreover, semantic communication systems may also need to be measured under such a quality consideration. Previous efforts in the study of coding with perception quality consideration have focused on the source coding formulation, i.e., the rate-distortion-perception function. Several conclusive results have been reported [5, 6, 7, 8, 9].
It should be noted that the perception quality consideration is different from the classical single-letter additive distortion measure, since the quality may be measured on the whole coding block instead of on individual sample positions. Therefore, the question of whether separation is optimal under the perception quality consideration remains mostly open. In this work, we revisit source-channel separation under the distortion-perception consideration. We show that when the perception quality is measured on the block level, i.e., in the strong sense, the optimality of separation still holds when common randomness is shared between the encoder and the decoder; however, separation is no longer optimal in general when such common randomness is not available. In contrast, when the perception quality is the average per-symbol measure, i.e., in the weak sense, the optimality of separation holds regardless of the availability of the common randomness.
Our result for the case under strong-sense perception measure with common randomness is particularly notable, since the proof does not rely on a single-letter characterization of the corresponding rate-distortion-perception region. This proof allows us to establish the separation result under a general set of conditions. We note that in a closely related work [10], the problem of source-channel separations was considered when the reconstruction needs to follow a fixed target distribution in an i.i.d. manner. In contrast, in this work, we consider distortion-perception coding, where the reconstruction does not need to satisfy any specific target distribution, nor does it need to be done in an i.i.d. manner.
@@ II-A Notation and Preliminaries
The channel we consider in this work is the classic discrete memoryless channel given by $P_{Y|X}$, and we denote its capacity as $C$. We assume the discrete memoryless source follows a probability distribution $P_{S}$ on a finite alphabet $\mathcal{S}$. A distortion measure function $\delta:\mathcal{S}\times\mathcal{S}\rightarrow[0,\infty)$ and a sequence of perception measure functions $d_{n}:\Delta_{\mathcal{S}^{n}}\times\Delta_{\mathcal{S}^{n}}\rightarrow[0,\infty]$ are assumed which satisfy $d_{n}(P,P)=0$, where $\Delta$ is the probability simplex on the alphabet specified by the subscript.
@@
We state several assumptions on the perception measure below, which are not always required for our results.
@@
Assumption 2.
Assumption 3.
The perception metric $d_{1}$ is convex in the second argument.
Assumption 1 allows us to establish separation results when single-letter characterizations are unavailable, and it is more lenient than the assumption in [10]. Several well-known measures, such as the squared Wasserstein-2 distance and the total variation distance, indeed satisfy these assumptions.
@@ II-B DP Joint Source-Channel Coding Achievable Regions
A distortion and perception pair $(D,P)$ is called achievable on a channel with a mismatch factor $\kappa$, if for all sufficiently large $n$, there exist common randomness $U$, an encoding function $f:\mathcal{S}^{n}\times\mathcal{U}\rightarrow\mathcal{X}^{m}$ (where $\mathcal{U}$ is a Polish space), a decoding function $g:\mathcal{Y}^{m}\times\mathcal{U}\rightarrow\mathcal{S}^{n}$, such that $\hat{S}^{n}=g(Y^{m},U)$ is the reconstruction and $Y^{m}$ is the channel output on the memoryless channel $P_{Y|X}$ with input vector $X^{m}=f(S^{n},U)$, which satisfy the condition that
@@
The first condition states that the source-channel mismatch factor must be upper-bounded properly, and the second and the last conditions specify the distortion constraint and the (strong-sense) perception constraint, respectively. There are related concepts in the literature that are also of practical interest. We consider the following cases in our work:
@@ II-C RDP Coding Achievable Regions
In the definition above, $Z$ can be viewed as an intermediate encoding of $S^{n}$ before the entropy coding. Similar to the joint coding settings, we can define the weak-sense rate-distortion-perception achievable region, the strong-sense rate-distortion-perception achievable region without common randomness, and the weak-sense rate-distortion-perception achievable region without common randomness, respectively. We denote these regions as $\mathcal{G}_{w}$, $\hat{\mathcal{G}}$, $\hat{\mathcal{G}}_{w}$, respectively.
@@
The distortion-perception joint coding achievable region $\mathcal{J}$ and the rate-distortion-perception achievable region $\mathcal{G}$ are both convex, regardless of whether the strong-sense or the weak-sense perception measure is adopted.
@@ II-D DP Separate Coding Regions
We can similarly define the weak-sense separate coding achievable distortion-perception region, the strong-sense separate coding achievable distortion-perception region without common randomness, and the weak-sense separate coding achievable distortion-perception region without common randomness, denoted as $\mathcal{J}^{\prime}_{w}$, $\hat{\mathcal{J}}^{\prime}$, and $\hat{\mathcal{J}}^{\prime}_{w}$, respectively.
We also need the following distortion-rate-perception function in the proof:
$D(R,P)=\min_{(R,D,P)\in\mathcal{G}}D.$ (10)
@@
III Coding under the Strong-Sense Perception Requirement
@@ III-A Separation: Common Randomness Present
$\mathcal{J}^{\prime}\supseteq\mathcal{J}$.
@@
Under Assumptions 1 and 2, we have $\mathcal{J}^{\prime}\subseteq\mathcal{J}$.
As a consequence, under Assumptions 1 and 2, we have $\mathcal{J}^{\prime}=\mathcal{J}$, which essentially says the separate coding achievable region is exactly the same as the joint coding achievable region under these conditions. We remark that a single-letter characterization of the achievable rate-distortion-perception region is not known under Assumptions 1 and 2. The single-letter results given in [7] were in fact established under a more stringent requirement on the perception functions. Nevertheless, we are still able to establish the optimality of the separation architecture using a proof that does not rely on the single-letter characterization.
@@ III-B Failure of Separation: Common Randomness Absent
Under the strong-sense perception requirement, when there is no common randomness shared by the encoder and the decoder, source-channel separation is not optimal in general. We next provide a counter-example from [10] to illustrate this.
@@ IV Coding under the Weak-Sense Perception Requirement
Under Assumptions 2 (which only needs to hold with $n=1$) and 3, $\mathcal{J}_{w}=\mathcal{J}^{\prime}_{w}=\hat{\mathcal{J}}_{w}=\hat{\mathcal{J}}^{\prime}_{w}=\mathcal{J}^{*}_{w}$, where
@@ V-A Proof of Theorem 1
We wish to show that there exists a rate-distortion-perception code that achieves with $R\leq\kappa C+\delta$ for any $\delta>0$. By the strong functional representation lemma [11], there exists a random variable $U^{\prime}$, independent of $S^{n}$ such that $\hat{S}^{n}$ can be expressed as a deterministic function of $(S^{n},U^{\prime})$ and
@@
for some positive $\delta_{n,m}$ which diminishes as $n,m$ become large. Since our new code preserves $p_{S^{n},\hat{S}^{n}}$, we have
@@ V-B Proof of Theorem 2
$\mathcal{J}\supseteq\{(D,P):(R,D,P)\in\mathcal{G},R<\kappa C\},$ (19)
instead of $\mathcal{J}\supseteq\mathcal{J}^{\prime}$ directly, because the joint coding achievable distortion-perception region $\mathcal{J}$ is closed by definition, and $D(R,P)$ is convex for any fixed value $P\geq 0$, and thus continuous in the relative interior.
@@
With this code, let us consider its length-$k$ extension, with $U_{i}$’s mutually independent for $i=1,2,\ldots,k$, which would allow us to address the event when the description $Z$ (conditioned on $U$) is too long to fit in a channel coding block. Let $P_{S^{nk},\hat{S}^{nk}}$ denote the joint distribution of $(S^{nk},\hat{S}^{nk})$ induced by this extended code.
When $k$ is large, we have $\frac{1}{nk}\sum_{i=1}^{k}H(Z_{i}|U_{i})\leq R+\delta$ holds with high probability. More precisely, choose an arbitrary pair $\delta_{R},\epsilon_{R}>0$, there exists a sufficiently large $k$ such that
$Pr(\text{the sum of description lengths of $Z_{1},\ldots,Z_{k}$}$
$\qquad\qquad\qquad\qquad\geq nk(R+\epsilon+\delta_{R}))\leq\epsilon_{R}.$ (23)
We shall send the cumulative descriptions of the $k$-extension of the original rate-distortion perception code, if the cummulative description length is less than $nk(R+\epsilon+\delta_{R})$, using an almost capacity-achieving channel code; otherwise, we simply send a fixed dummy sequence of a shorter length (which will be viewed as an error event correspondingly).
With $\delta_{R},\epsilon$ chosen to be any fixed small quantities, since $R<\kappa C$, when $n,k$ are both sufficiently large, there exists a channel code at rate greater than $\kappa^{-1}(R+\epsilon+\delta_{R}+\frac{1}{nk})$, such that the decoding error probability is upper-bounded by $\epsilon_{C}$, where $\epsilon_{C}$ diminishes as $n,k$ are sufficiently large. The new decoder first attempts to recover the intended $Z^{k}$ sequences, and let $\hat{Z}^{k}$ denote this reconstruction. Two error events (when $Z^{k}\neq\hat{Z}^{k}$) may occur: 1) The description of the concatenated source sequence $Z^{k}$ is too long and the fixed dummy sequence was sent, and 2) A channel decoder error occurred. By the union bound, we have
@@
$\leq\frac{1}{nk}\sum_{t=1}^{nk}\mathbb{E}_{P}[\delta(S_{t},\hat{S}_{t})]+2(\epsilon_{R}+\epsilon_{C})D_{\max}$
$\leq D+\epsilon+2(\epsilon_{R}+\epsilon_{C})D_{\max},$ (26)
where we have used eqn. (29) in [13], and
@@
where (27) is due to the fact that $d_{nk}$ is continuous with respect to $d_{TV}$ (i.e., Assumption 2) and sub-decomposable. The code clearly satisfies the channel mismatch factor $\kappa$. ∎
@@ V-C Proof of Theorem 3
In the weak-sense setting, it is clear by definition that $\mathcal{J}_{w}\supseteq\hat{\mathcal{J}}_{w}$, $\mathcal{J}_{w}\supseteq\mathcal{J}^{\prime}_{w}$, and $\hat{\mathcal{J}}_{w}\supseteq\hat{\mathcal{J}}^{\prime}_{w}$. It was also shown in [7] that under Assumption 2 (which only needs to hold for $n=1$) and Assumption 3, the weak-sense achievable rate-distortion-perception region is characterized as
@@
To show $\mathcal{J}_{w}\subseteq\mathcal{J}^{*}_{w}$, consider an arbitrary pair $(D,P)\in\mathcal{J}_{w}$. For any $\epsilon>0$, there exists a perception-distortion code $(f,g)$ that achieves $(D+\epsilon,P+\epsilon)$. For this code, we have
@@
$P+\epsilon\geq\frac{1}{n}\sum_{t}\mathbb{E}d_{1}(S_{t},\hat{S}_{t})\geq\mathbb{E}d_{1}(S,\hat{S}),$ (33)
where Assumption 3 is used in (33). Since $\epsilon$ can be made arbitrarily small by making $n$ large, we have $(D,P)\in\mathcal{J}^{*}_{w}$.
To show $\hat{\mathcal{J}}^{\prime}_{w}\supseteq\mathcal{J}^{*}_{w}$, we shall rely on the combination of the coding scheme given in [7] for the rate-distortion-perception problem and the conventional capacity-achieve channel codes. Similarly to the proof of Theorem 2, the only issue that needs to be resolved is that due to excess source coding description length and channel uncertainty, the description $Z$ has a diminishing error probability of being decoded erroneously at the decoder. However, the penalties such induced on the distortion and the perception are both negligible, when Assumption 2 holds with $n=1$. The proof is thus complete. ∎
@@ VI Conclusion
We consider source-channel separation with perception consideration in addition to the standard distortion measure. We show that under the strong-sense perception constraint, separation is optimal when common randomness is available at the encoder and the decoder; however, when common randomness is not available, source-channel separation is not optimal in general. In contrast, under the weak-sense perception constraint, separation is always optimal regardless of the existence of common randomness.
Acknowledgement
The authors are grateful to the anonymous reviewer who provided numerous constructive comments, which helped to improve the quality of this work.